### v1

# The Past, Present, and Future of Automation in Model-Driven Engineering

Abstract.
Model-Driven Engineering (MDE) provides a huge body of knowledge of automation for many different engineering tasks, especially those involving transitioning from design to implementation. With the huge progress made on Artificial Intelligence (AI) techniques, questions arise for the future of MDE such as how existing MDE techniques and technologies can be improved or how other activities which currently lack dedicated support can also be automated. However, at the same time, it has to be revisited where and how models should be used to keep the engineers in the loop for creating, operating, and maintaining complex systems. To trigger dedicated research on these open points, we discuss the history of automation in MDE and present perspectives on how automation in MDE can be further improved and which obstacles have to be overcome in the medium and long term perspective.
conference: International Workshop on Software Engineering in 2030; July 2024; doi:
1. Introduction
Bézivin coined the term Model-Driven Engineering (MDE) in 2005 as follows: “MDE is a software engineering methodology which uses formal models, i.e., models which are machine-readable and processable, to semi-automatically produce executable software systems” (Bézivin, 2005). The main idea behind formal models is to provide abstractions of software systems to manage their complexity. Based on these abstractions, as they are machine-readable and even processable, automation techniques are provided, e.g., to generate executable software (Schmidt, 2006), which was the main focus 20 years ago in the pioneering phase of MDE. However, soon, other automation concerns emerged and have been studied, such as analysing models by transforming them into formal domains such as model checkers. Model-to-model transformations have been used for this purpose. For code generation, model-to-code transformations have been developed to automate the transition from design models to code-based implementations. To complete the picture, code-to-model transformations have been used for reverse engineering of existing software systems such as legacy systems. A comprehensive study on different transformation intents is provided in (Lúcio et al., 2016).
For developing automation support in terms of transformations, major research efforts have been spent on developing model transformation languages as it has been soon realized that transformations are complex software systems themselves (Sendall and Kozaczynski, 2003). In parallel, language engineering approaches based on metamodeling techniques have been emerging, which allow for the efficient and effective development of modeling languages and accompanying modeling tool support such as model editors, validators, and storage facilities (Atkinson and Kühne, 2003). Having these infrastructures enabled the development of domain-specific modeling languages (DSMLs) (Chen et al., 2005) in addition to general-purpose modeling languages such as UML.
By having abstractions as models combined with automation support, the promise of MDE is to allow engineers to focus on the problem domain, i.e., which systems should be realized, rather than on the solution domain, i.e., how the system is realized. However, to realize this promise, several challenges have to be tackled as reported over the years in several papers. In (France and Rumpe, 2007), the authors discuss challenges associated with modeling languages, e.g., how to engineer DSMLs, achieving separation of concerns to deal with complexity, and model manipulation/management mostly supported by transformations. In (Mussbacher et al., 2014), the authors discuss how MDE may be used even beyond software, raising the need for interdisciplinary modeling support, e.g., for socio-technical systems. The latter requires model integration and consistency support as well as strong tool support. Finally, in a more recent work (Bucchiarone et al., 2020), the authors discuss important aspects of MDE such as foundational, social, community, domain, and tool aspects. However, we are not aware of any report focusing explicitly on the automation aspect of MDE and how this aspect will develop towards 2030. It is important to close this gap as with the recent and expected advances of AI, the current automation techniques and technologies of MDE are highly impacted as well and new ones may be developed to support additional tasks in the software engineering process which are not targeted by existing technologies. Thus, a discussion on the role of models, MDE methodologies, and engineers in this new era is required to frame the potential future of MDE.
Important questions have to be answered, such as which phases can be additionally automated, e.g., should we directly go from natural language requirements to code, or is there still, especially because of the enhanced AI capabilities, a sweet spot for using models in such highly automated processes? Can we now ease and automate the modeling process itself and reach at the same time a higher automation degree in the synthesis process, such as code generation? These questions are important for the foundations of MDE, but at the same time, also for recent application areas such as low-code/no-code development (Ruscio et al., 2022) and digital twin engineering (Tao et al., 2018).
In the following, we capture the achievements of automation in MDE by enumerating prominent engineering activities which have been in the focus of automation as well as the automation techniques which have been applied for achieving the automation. Based on this foundation, we lay out several challenges hindering further significant automation increases and conclude the paper with perspectives on tackling these challenges until 2030.
2. Achievements in Automation
2.1. Targeted Automation Activities
In this section, we categorize and discuss the activities that have been targeted for automation within the MDE field, organizing them according to the categories outlined in the contents for a Model-Based Software Engineering Body of Knowledge (Burgueño et al., 2019).
2.1.1. Model Quality Management
Significant efforts have been made to automate activities for finding defects, faults and failures in models and modeling processes. Numerous model validation and verification techniques exist that automate quality checks, such as correct syntax, semantics, structure, behaviour, etc., of models to ensure that any inconsistencies or errors are detected and rectified at the model level and as soon as possible. For instance, model checking has been automated through algorithms that check the temporal behaviour of a software or system model against certain properties or specifications, ensuring they meet specified requirements or constraints, and there are numerous languages and mature tools available (Holzmann, 2004; Bengtsson et al., 1995; Kwiatkowska et al., 2011).
Automated testing techniques based on models (a.k.a. model-based testing) have been developed, too, where models are used to generate test cases automatically (e.g., (Pretschner et al., 2005)). Traditionally, this method has helped identify potential problems early in the development process, significantly reducing cost and effort in the testing phase. Moreover, in recent years, model-based testing automation has gained relevance due to the rise of digital twins (Bickford et al., 2020), which impose the need for streamlining testing, and the adoption of new development methodologies such as agile, continuous development, integration and delivery (Jongeling et al., 2019) where the process of executing test cases and analyzing the results, often in a continuous integration / continuous deployment (CI/CD) pipeline that needs to be automated.
2.1.2. Model-based execution
For testing, debugging or analysis, executable models and simulation facilities are necessary. Regarding executable models (Kleppe et al., 2003), executable UML models are popular, which usually use an action language to define the behaviour of the system (e.g., the Alf Action Language (alf, 2017), SOIL (Schäfer and Gogolla, 2020), or FUML [^1] https://www.omg.org/spec/FUML).
Over the years, the automation of model-based execution, simulation, and co-simulation has seen numerous advancements, moving from manual processes that required human intervention—e.g., to set up simulations and adjust parameters based on expert knowledge and observations—to highly automated processes and tools. For instance, there has been the standardization of an Action Language as part of the OMG UML Specification for the execution of UML models (alf, 2017) and tools for simulation such as Modelica [^2] https://modelica.org, which are widely used by system engineers. There is also available a series of interpreters for modeling languages, both graphical and textual, many of them fully integrated into modeling environments, e.g., see (Gogolla et al., 2007).
[…]
2.1.3. Model Transformations
Model transformations constitute a fundamental aspect of MDE and serve as the backbone for automating the transition from high-level abstractions to executable code, simulations, or other models. They can be classified into Model-to-model transformations (M2M), Model-to-text (M2T) transformations, and Text-to-model (T2M) transformations (Kleppe et al., 2003). M2M transformations are perhaps the most exploited automation and automated activity in MDE, where, among other uses, models are automatically transformed into executable code in various programming languages. M2T transformations provide the ability to autogenerate code from models, contributing significantly to software development automation. The models serve as the primary artifact; from them, code, documentation or other textual artifacts (such as JSON files) are generated automatically (Klein, 2015). T2M transformations are used for reverse-engineering code into models as well as other purposes such as documentation generation (Wang et al., 2010).
At the same time that model transformations have enabled and pushed forward automation in software engineering, they have also been the target of automation. This automation leverages specific languages, execution engines, tools, and frameworks, as described below. One of the primary means of automating model transformations is through dedicated transformation languages. These languages provide constructs tailored to describe how model elements must be mapped to their equivalent elements. Examples of such languages include ATL (Jouault et al., 2008), QVT (OMG, 2005), ETL (Kolovos et al., 2008), and Henshin (Arendt et al., 2010) to name just a few—for a survey see (Kahani et al., 2019).
Transformation engines [^4] https://projects.eclipse.org/projects/modeling.mmt.qvt-oml [^5] https://wiki.eclipse.org/ATL/EMFTVM(Burgueño et al., 2016; Cuadrado et al., 2022; Tisi et al., 2013) are specialized software components that execute the transformation rules defined in a transformation language against source models to generate target models. These engines manage the complexities of navigating model structures, applying transformation rules, resolving dependencies, and ensuring the integrity of the generated models. Transformation engines can be standalone tools or integrated into broader MDE platforms, providing a seamless experience from model editing through to transformation.
For automation to be truly effective, model transformations must be well-integrated with the software development environment. This includes providing user-friendly interfaces for defining and managing transformations. Such integration enables developers to incorporate model transformations seamlessly into their workflows, automating repetitive tasks, and facilitating more agile and iterative development processes. Some efforts include the Epsilon Playground (Kolovos and García-Domínguez, 2022) and AToMPM (Syriani et al., 2013) among others.
2.1.4. Model Management
Model management involves different activities, which we discuss in the following. Automated model evolution involves strategies for evolving models to reflect changes in requirements or design. Automation can be achieved through predefined transformation rules or templates that apply changes across models, ensuring consistency and reducing manual effort. Languages such as those that conform to Epsilon (Kolovos and García-Domínguez, 2022) or ATL (Jouault et al., 2008) support automated transformations that facilitate model evolution.
When metamodels evolve, models that conform to those metamodels may become invalid and need to be adapted to remain consistent with the new version of the metamodel (Herrmannsdoerfer and Ratiu, 2009). Automated model migration is needed to update models. So far, this has been achieved in numerous ways, for instance, by means of model transformations (Sprinkle and Karsai, 2004; Ruscio et al., 2012) or with dedicated languages such as Epsilon Flock (Rose et al., 2014).
Maintaining models to ensure their accuracy and relevance over time can be automated through continuous integration tools that integrate model changes, run automated validations, and report inconsistencies or errors. Automated testing frameworks for models also contribute to maintenance by ensuring that model updates do not break existing functionalities.
Since models are software artifacts, version control systems are beneficial to keep up with current development needs. Given the different nature of models w.r.t. code, approaches to deal with models are needed. For instance, EMFStore (Koegel and Helming, 2010) or Connected Data Objects (CDO) [^6] https://www.eclipse.org/cdo provide automated support for model versioning.
Automated model diffing tools compare and identify differences between model versions. EMF Compare [^7] https://eclipse.dev/emf/compare is an example of a tool that provides automated support for model comparison, highlighting differences and conflicts within models developed using EMF (Eclipse Modeling Framework) [^8] https://eclipse.dev/modeling/emf.
Merging model changes, especially in collaborative environments or when reconciling parallel development efforts, require sophisticated automation to handle conflicts and ensure model integrity. Tools like EMF Compare [^9] https://eclipse.dev/emf/compare also support automated model merging by providing strategies and user interfaces to resolve conflicts and integrate changes from different model versions seamlessly.
Refactoring models is the process of restructuring the model while not modifying its original functionality or behaviour. This process can be automated using model transformation languages and refactoring libraries. Tools and frameworks, such as Epsilon (Kolovos and García-Domínguez, 2022) or EMF Refactor (Arendt and Taentzer, 2012),provide automated support for model refactoring.
2.1.5. Modeling
Last but not least, the task of modeling itself, i.e., creating models as abstractions of software and systems, has also been the target of automation. In particular, specific efforts have been put into approaches for model generation (Saini et al., 2022), model search (López and Cuadrado, 2020), model completion (Chaaben et al., 2023) and model classification (Nguyen et al., 2019). Among other efforts, a long-lasting line of work to automate the execution of models has been heavily addressed by researchers working on models@runtime (Bencomo et al., 2019), where the models of a system need to be up-to-date while the system is running at the same time that is used to adapt the runtime behaviour.
2.2. Automation-Enabling Techniques
From the very beginnings of MDE, automation emerged as a fundamental pillar. Numerous works have employed direct approaches based on heuristics to automate or semi-automate certain activities directly on the models’ representations. Other works have relied on established automation technologies to achieve this automation. In the remainder of this section, we discuss four families of these technologies: formal methods, search-based techniques, extensional and intensional knowledge engineering approaches, and artificial intelligence.
2.2.1. Formal Methods
Due to their rigor and precision, formal methods have been utilized for specifying, analyzing, and verifying various artifacts in MDE. The literature is abundant with contributions on the application of formal methods for automating various MDE tasks. However, we limit ourselves to a few examples of their most common uses.
The first family of applications targets the analysis and verification of UML models. This involves techniques such as model checking (Hilken et al., 2018), constraint satisfaction problem solving (Ehrig et al., 2009), Boolean satisfiability (Soeken et al., 2010), and rewriting logic (Durán, 2022). Another family of application of formal methods concerns the generation of models for metamodel or transformation testing. Here again, a variety of techniques have been explored, including typed graph transformation (Anastasakis et al., 2010), equivalence partitioning (Cabot et al., 2007), and satisfiability modulo theories (Semeráth et al., 2020), to name a few. A third noteworthy family, although few works have covered it, is model completion. Its automation is achieved, for example, through first-order logic using Alloy (Sen et al., 2010). Formal methods have also been applied to define the operational semantics of modeling languages, e.g., by using Maude and graph transformations (Rivera et al., 2009a; Engels et al., 2000). Formal methods have also been employed for model versioning (Westfechtel, 2014; Taentzer et al., 2014), including diffing and merging, which have been managed in terms of graph modifications and graph transformations.
2.2.2. Search-based Techniques
These techniques have emerged as valuable tools for automating various activities of MDE. This line of research views the automation of many MDE tasks as optimization problems, which can be effectively tackled using search-based methods (Boussaïd et al., 2017). One of the earliest applications of this concept was in the domain of model transformations, particularly using examples of input-output models. Researchers utilized search-based techniques for model diffing based on refactoring detection(Kessentini et al., 2017), model merging (Mansoor et al., 2015), and to perform model transformations (Kessentini et al., 2008) and even to learn transformation rules (Faunes et al., 2013; Baki and Sahraoui, 2016).
Moreover, search-based techniques have been instrumental in generating test suites to validate these transformations. Test cases, often in the form of input-output models, are generated to ensure the correctness of the transformations (Jilani et al., 2014). Additionally, model samples play a crucial role in testing metamodels. Search-based techniques have been employed to either generate those model samples (Gómez et al., 2012) or to perform model selection (Batot and Sahraoui, 2016) to ensure the robustness and effectiveness of metamodels. In cases where metamodel descriptions lack well-formedness rules, these rules can be recovered from examples of both valid and invalid models. Genetic programming techniques have been utilized for this purpose, aiding in the refinement of metamodels (Batot and Sahraoui, 2022).
Another significant application of search-based techniques in MDE is the repair of modeling artifacts. The evolution of metamodels can invalidate many associated artifacts due to their dependency on the type systems defined within the metamodels. Evolutionary algorithms have been employed to repair models (Kessentini and Alizadeh, 2022), OCL constraints (Batot et al., 2017), and model transformations (Kessentini et al., 2018a). Similarly and independently from the evolution, search-based approaches have been used to repair errors in MDE artifacts such as transformations (VaraminyBahnemiry et al., 2021).
In the realm of design space exploration (DSE), where finding the best design instance for a given problem is challenging, search-based techniques have proven invaluable. DSE, framed as an optimization problem, involves exploring alternative models to identify those that meet specific criteria (Abdeen et al., 2014).
The relation between MDE and search-based techniques is also a tale of cross-fertilization. Indeed, MDE contributed to enriching search techniques through improved encoding of optimization problems and their subsequent resolution, e.g.,(Denil et al., 2014; Burdusel et al., 2018; John et al., 2023; Bill et al., 2019).
2.2.3. Extensional and Intensional Knowledge
The automation or semi-automation of a task requires knowledge that can be operationalized through software tools. This knowledge may include elements that enable the direct execution of the task. It can also be exploited within a complex automation process. In the context of MDE, knowledge often appears extensionally in some kind of repositories, particularly in the form of sets of artifacts such as models (Robles et al., 2017) or transformations [^10] https://eclipse.dev/atl/atlTransformations. It can also be intentionally represented, typically through ontologies.
Among the artifacts within MDE, model sets have been extensively leveraged to provide support in modeling tasks. These collections of models are employed via search engines to propose models or model fragments aligning with specific criteria articulated by modelers, either in textual or pre-defined partial model forms (Bislimovska et al., 2014; López and Cuadrado, 2022; Basciani et al., 2018a).
The alternative category of knowledge, termed intensional, integrated into MDE task automation, encompasses ontologies. These structures provide organized descriptions of real-world objects, which can be utilized for crafting domain-specific languages or models (Staab et al., 2010). Due to their inherent semantics, ontologies also serve in model verification and debugging processes (Mokos et al., 2010; Walter et al., 2014), or for model matching (Kappel et al., 2006). Additionally, ontologies can be harnessed through design patterns to automate the creation of meta-models or models (Atkinson and Kühne, 2003; Pescador and de Lara, 2016).
Lastly, the integration of extensional and intensional knowledge can enhance automation efforts. For instance, ontologies can streamline search requests and responses within model repositories (Robles et al., 2012).
2.2.4. Artificial Intelligence
Like many other disciplines, AI has been extensively used as an automation technology for MDE. From the early days of MDE until now, rule-based systems have been employed to automate model transformations, utilizing either general-purpose rule languages such as NéOpus (Revault et al., 1995) or specifically dedicated ones like ATL (Jouault et al., 2008). Soon after, the notion of supporting modelers in writing transformation rules emerged. In this regard, machine learning techniques were applied to derive transformation rules from examples using supervised learning (Varró and Balogh, 2007) or unsupervised learning (Dolques et al., 2009).
More recently, with advancements in machine learning research, considerable effort has been dedicated to assisting in metamodeling and modeling activities. This includes the utilization of techniques such as GNNs (Di Rocco et al., 2021) and LSTM (Weyssow et al., 2022). Furthermore, research has been focused on model transformations using reinforcement learning (Eisenberg et al., 2021) and LSTM architectures (Burgueño et al., 2022). Various machine learning algorithms have also been applied for metamodel classification (Nguyen et al., 2019) and model classification (López et al., 2022), as well as for model repair (Barriga et al., 2022) and model generation (López et al., 2022). However, these initiatives often face challenges due to the difficulty of training machine learning models, particularly given the scarcity of data.
The remarkable power of LLMs has partially alleviated this problem. Thus, it became possible to automate several modeling tasks with little or no training data. However, the effectiveness of using LLMs for this purpose varies, as Camara et al. (Cámara et al., 2023) highlight. Ben Chaaben et al. explored the application of LLMs for static and dynamic domain-model completion (Chaaben et al., 2023). Model generation with LLMs has also garnered significant interest among researchers, encompassing diverse model types such as goal models (Chen et al., 2023a) and static domain models (Chen et al., 2023b).
2.3. Mapping between Activities and Enabling Technologies
The aim of this subsection is to present a mapping between the previously presented modeling activities and the enabling technologies. This mapping does not intend to be exhaustive by any means but wants to present illustrative examples. Table 1111111Since a metamodel is a model (in particular, a model of a model), some of these activities apply to both models and metamodels. For the sake of brevity, we only refer to models. shows activities in its rows and technologies in its columns (the column “Direct” contains examples of approaches providing automation support directly for the model representations without moving to an already existing automation space). Each cell contains a maximum of two exemplary works, appended with a ‘+’ if the state of the art is more extensive and includes further works. Cells with only one work mean that we could not find more examples for that activity and technology. Empty cells mean that we could not find any example.
Activities Enabling Technologies
Direct Formal Methods Know. Eng. Traditional AI* Search Techniques Deep Learning** LLMs
Model quality management Analysis / Verification (Basciani et al., 2019; Ma et al., 2013)+ (Gogolla et al., 2018; Rivera et al., 2009a)+ (Mokos et al., 2010) (Barriga et al., 2022; Wenz et al., 2021)+
Testing / Debugging (Popoola et al., 2016; Bagherzadeh et al., 2017)+ (Hilken et al., 2018; Babikian et al., 2022)+ (Walter et al., 2014) (Batot and Sahraoui, 2016; Gómez et al., 2012)+ (Rahimi et al., 2023; López and Cuadrado, 2023)+
Model transformations (Cuadrado et al., 2022; Biermann et al., 2006)+ (Biermann et al., 2008; Gammaitoni and Kelsen, 2019)+ (Roser and Bauer, 2005; Kappel et al., 2006)+ (Eisenberg et al., 2021) (Kessentini et al., 2008; Baki and Sahraoui, 2016)+ (Burgueño et al., 2022; Lano and Xue, 2023)
Model based execution (Muller et al., 2005; Combemale et al., 2017)+ (Rivera et al., 2009a; Engels et al., 2000)+
Model management
Model (co-)evolution
(Ruscio et al., 2012; Herrmannsdoerfer, 2011)+ (Taentzer et al., 2012; Levendovszky et al., 2014)+ (Kessentini et al., 2018b; Kessentini and Alizadeh, 2022)+
Model migration and maintenance
(Rose et al., 2014; Herrmannsdoerfer and Ratiu, 2009)+ (Giese and Wagner, 2006) (Barriga et al., 2020)
Model versioning / diff / merge
(Koegel et al., 2010; Brun and Pierantonio, 2008)+ (Westfechtel, 2014; Taentzer et al., 2014)+ (Kappel et al., 2006) (Kessentini et al., 2017; Mansoor et al., 2015)+
Model refactoring
(Sunyé et al., 2001; Porres, 2003)+ (Mens, 2005; Biermann et al., 2006)+ (Mokaddem et al., 2018; Ghannem et al., 2018)+
Modeling
Model generation
(Gómez-Abajo et al., 2016; Brottier et al., 2006)+ (Babikian et al., 2020; Schneider et al., 2017)+ (López et al., 2022) (Jilani et al., 2014; Abdeen et al., 2014)+ (Chen et al., 2023b, a)
Model search
(Lucrédio et al., 2008; Basciani et al., 2018b)+
Model completion
(Kuschke et al., 2013; Born et al., 2008)+ (Sen et al., 2010; Steimann and Ulke, 2013)+ (Bislimovska et al., 2014; López and Cuadrado, 2022)+ (Nechypurenko et al., 2007) (Di Rocco et al., 2021) (Chaaben et al., 2023)
Models at runtime
(Bencomo et al., 2008; Fouquet et al., 2012)+ (Búr et al., 2018; Sakizloglou et al., 2022)+ (Bencomo et al., 2013; Hartmann et al., 2019)+ (Kinneer et al., 2021; Eisenberg et al., 2022)
Model classification
(Rubei et al., 2021; Basciani et al., 2016)+ (Nguyen et al., 2019) (Nguyen et al., 2021)
    Excluding Search-based Techniques
    Excluding LLMs
Table 1. Overview of MDE activities and enabling technologies.
Table 1121212The table is inevitably nonexhaustive and merely intended to reflect the kinds of activities and supporting approaches available. gives a partial summary of how automation-enabling technologies have been used for various MDE tasks. As we can see, a range of techniques, including direct programming, formal methods, and search methodologies have been extensively employed to automate various MDE tasks. To a lesser extent, the MDE community have leveraged, on the one hand, model sets and ontologies and, on the other hand, other AI techniques to automate some distinct activities. On another note, in recent years, there has been a surge in interest and adoption of deep learning-based approaches for MDE, albeit their widespread use has been hindered by the scarcity of large-scale datasets for model training or refinement.
[…]
To capitalize on these opportunities, there is a pressing need to prioritize the creation of expansive datasets tailored specifically for MDE tasks, as highlighted in Section 3. By doing so, researchers and practitioners can unlock the full potential of deep learning methodologies in automating and enhancing various tasks of MDE.
Looking ahead, LLMs are poised to become a pivotal research domain in automating MDE tasks. However, their application extends beyond mere model generation or completion tasks. There exists a vast scope for creative exploration and utilization of LLMs in diverse MDE activities, heralding a new era of innovation and efficiency in software engineering practices.
3. Perspectives and Challenges
Building upon the introductory concepts discussed earlier, this section provides an overview of perspective topics that hold significance in the medium to long term for automation in MDE. We foresee a vast potential for automation within MDE, driven by advancements in AI, particularly in generative AI and LLMs. In exploring the future prospects of MDE and its automation possibilities, it is crucial to acknowledge the challenges ahead. Challenges include integrating automation tools, ensuring compatibility, and maintaining interoperability among diverse modeling languages and environments. Additionally, developing intuitive modeling tools and addressing ethical considerations in interdisciplinary domains pose significant problems to be investigated, as discussed below. The distinction between perspectives and challenges may not always be clear-cut and may overlap. While discussing envisioned advancements, we see some of them as potential obstacles for automation that must be overcome.
Management of ever-increasing diversity of languages and tools: MDE has seen a proliferation of domain-specific modeling languages and tools, which has led to a diverse range of technologies. However, this diversity has posed challenges for generic automation as each modeling language has its own syntax, semantics, and transformation rules. Despite efforts to standardize syntax, the MDE field still needs comprehensive standards, especially concerning semantics, transformation rules, and model interchange formats. Moreover, the rise of low-code/no-code platforms makes this issue even more difficult as adherence to existing standards becomes increasingly sporadic. Thus, it is imperative to develop standardized interfaces, transformation mechanisms, and model interchange formats to ensure interoperability and tooling support across different formalisms, languages, and paradigms.
Combination of symbolic and non-symbolic AI: Integrating symbolic AI techniques, such as logic reasoning and rule-based systems, with non-symbolic approaches, such as neural networks and deep learning, offers opportunities for enhancing automation in MDE. By leveraging the complementary strengths of both paradigms, researchers can develop hybrid AI systems that excel in tasks requiring both symbolic reasoning and pattern recognition. For example, combining symbolic reasoning with neural networks enables MDE tools to interpret and manipulate models based on logical constraints while leveraging the scalability and flexibility of neural networks for data-driven tasks or even exploiting generative AI tools to synthesize models that are domain or context-aware.
Emancipation from formal languages with LLMs: Creating domain-specific modeling languages is a crucial undertaking that involves collaboration between different roles, including language and domain experts. This collaborative effort encompasses various steps, including developing concrete syntaxes, the definition of formal semantics, and the construction of supportive modeling environments. However, the language development process can be challenging, often necessitating refinement of language definitions to refine user requirements or address unforeseen ones. LLMs offer a promising option for the early evaluation of languages by specifying their semantics in natural language and facilitating early analysis and stakeholder interactions to validate whether the language aligns with user requirements.
Template-based code generation vs. AI-based code generation: While template-based methods provide transparency and control over the generated code, AI-based approaches offer adaptability and potential for optimizing code generation based on diverse requirements and contexts. Researchers can develop code-generation techniques by exploring hybrid approaches that combine the strengths of both paradigms.
Model management automation with AI approaches: AI-driven techniques can enhance model management processes by improving semantic understanding and context-aware manipulation of models. By leveraging AI for tasks such as model differencing, merging, and versioning, MDE can achieve greater efficiency and reliability in managing evolving model artifacts. For example, AI algorithms can analyze the semantics of models to detect and resolve (or propose how to resolve) conflicts during model merging, thereby reducing manual effort and potential errors in model integration.
Moving from general LLMs to task-specific AI agents: In the traditional context of MDE, a platform typically denotes a particular technology stack or execution environment. However, with the increasing adoption of LLMs, the platform concept expands to encompass a collaborative ecosystem of multiple language models operating together within a multi-agent system. Specifically, we foresee a paradigm shift towards the widespread adoption of task-specific AI agents driven by LLMs. These specialized agents are designed to execute various model management tasks, such as model evolution, comparison, domain modeling, training and test data generation, and model completion (Hong et al., 2023).
Orchestration of automated MDE tasks: The application of modeling concepts and tools involves executing different tasks, including model transformations, validation, and code generation. To this end, despite the availability of domain-specific languages like Wires (Rivera et al., 2009b), there is still the need for higher-level automation strategies that, starting from user requirements, can coordinate low-level automated tasks and that are flexible to accommodate different modeling languages, tools, and environments. Additionally, achieving a seamless integration of automated processes requires addressing compatibility issues and ensuring interoperability between different automation tools and frameworks.
Democratizing the usage of task-specific AI agents: To enable the usage of task-specific AI agents, it is necessary to devise low-code platforms that empower users to leverage task-specific AI agents seamlessly. Such platforms can facilitate broader participation in MDE activities across domains and expertise levels by providing intuitive interfaces for coordinating diverse AI models. Moreover, low-code platforms can serve as orchestrators for deploying and managing AI agents, allowing users to customize and extend automation capabilities according to their specific requirements and preferences.
Combination of physical/biological/social/software aspects in complex systems: Integrating physical, biological, social, and software aspects into the same system presents a multifaceted challenge that requires a comprehensive approach to modeling and abstraction. MDE stands as an ideal integration vehicle for managing these heterogeneous aspects by providing a unified framework for representing and analyzing complex systems. Digital twins, which are virtual replicas of physical entities or systems, offer a powerful abstraction mechanism for simulating and analyzing the behavior of interconnected systems. By leveraging digital twins, MDE can facilitate the design and optimization of integrated systems across diverse domains, ranging from healthcare to sustainable smart cities.
Emergence of non-traditional interface devices: Integrating speech interfaces, gesture recognition, and other non-traditional input modalities in MDE tools can democratize software development by enabling a broader range of stakeholders to participate in modeling activities. By providing intuitive and accessible interfaces, such as voice commands or touch-based interactions, MDE tools can lower the barriers to entry for users with diverse backgrounds and abilities. Moreover, non-traditional interface devices facilitate collaboration and communication among distributed teams, allowing users to interact with models in real time and co-create solutions collaboratively. To this end, interdisciplinary work at the intersection of human-machine interaction and software engineering research fields will be needed (Kashfi et al., 2019).
Feedback-driven MDE: MDE tools and processes will need built-in feedback mechanisms. This feedback could come from engineers, stakeholders or the system itself (i.e., in the form of real-time data) and could be sporadic or continuous. Tools, processes and teams should not disregard feedback. New methodologies to work in a new environment where automated feedback integration are needed, too. A straightforward scenario is Digital Twins. For instance, Digital Twins can be used for product development, where design and engineering teams can investigate a broader range of design possibilities early in the development, saving money, resources and time. In such a scenario, the engineering processes must be adapted to collect and react to the outputs/feedback received by the Digital Twin, and engineering teams must have methodologies put in place.
Automated and semi-automated decision-making in MDE: As software and systems grow in complexity and the demands for faster delivery increase, automating decision-making processes becomes needed to improve efficiency and quality. On a purely technical level, automated decision-making will help choose the best transformation languages, model checkers, AI algorithms, execution models, etc. Furthermore, it would ensure that decisions are made consistently throughout the life of a project and by different teams. It would also prevent human error in repetitive decision-making processes (e.g., tools enhanced with automated support could exploit data and human-defined objectives to make or suggest optimal decisions).
Scarcity of datasets and benchmarks for AI models: The advancement of AI-driven automation in MDE is hindered by the scarcity of high-quality datasets and benchmarks for training and evaluating AI models. Unlike domains like natural language processing or computer vision, where large-scale datasets are readily available, MDE lacks comprehensive datasets that capture a diverse range of modeling tasks and application domains (López et al., 2024). Moreover, concerns regarding data privacy and the ethical use of data further complicate the acquisition and sharing of modeling datasets. Addressing these challenges requires collaboration between researchers, practitioners, and stakeholders to curate datasets, define evaluation metrics, and establish best practices for data collection and sharing in MDE contexts (Rocco et al., 2015).
The advent of LLMs presents new opportunities for dataset creation in various domains. With the ability to generate human-like text and simulate expert knowledge, LLMs offer a cost-effective alternative to traditional human labeling methods. For instance, LLMs can automatically generate labeled datasets for MDE tasks such as model classification or sentiment analysis. Additionally, LLMs can serve as judges to evaluate the quality and accuracy of tasks performed by other LLMs, thereby facilitating continuous improvement and refinement of AI-driven systems.
Better modeling tools: Despite significant advancements in modeling tools, many existing tools in MDE are tailored towards expert users and lack intuitive interfaces for non-specialists. This poses challenges for automation and democratizing the adoption of MDE methodologies across diverse domains and user communities (Ruscio et al., 2022). Improving the accessibility and usability of modeling tools requires incorporating user-centered design principles, conducting usability studies, and providing comprehensive documentation and training materials. Additionally, ensuring compatibility and interoperability between different modeling tools and environments is essential for facilitating seamless collaboration and exchange of models among users with varying expertise levels.
Teaching people how to prompt: Despite the advancements in AI and automation, human expertise remains indispensable in designing complex systems and formulating effective prompts for AI models. Teaching people how to prompt effectively involves imparting knowledge and skills in understanding system requirements, defining clear objectives, and structuring prompts to elicit desired responses from AI models. Thus, human expertise is essential for ensuring the accuracy, relevance, and ethical considerations of AI-generated outputs.
Teaching MDE: Given the rapid evolution of technology and MDE, teaching MDE needs to be adapted to the fast-evolving needs. Educators must prepare students not just to be proficient modelers, but to be architects of highly complex software and systems that leverage automation and MDE techniques to meet the demands of industry and society. Incorporating new technologies in teaching MDE can significantly improve the learning experience. For instance, the use of simulation tools, multidimensional visualization software, and other interactive applications can provide a more immersive and engaging learning experience.
(Self-)Adaptation and coexistence with physical components: Using abstractions to automate the generation and adaptation of software will narrow the gap between the digital and physical worlds. This will have an important application to domains such as IoT, CPS or Digital Twins. Technical, security, and standardization objectives need to be defined and pursued to make MDE up to the current needs.
4. Conclusion
MDE has proven its utility over time, primarily due to its remarkable capabilities in abstraction and automation. These features have significantly streamlined the development process, from conceptualizing ideas to their implementation. This article delves into the various facets of MDE activities and the diverse technologies that have been deployed to automate these activities.
The evolution of MDE is closely intertwined with the emergence of cutting-edge technologies, such as deep learning and large language models. These advancements have revolutionized automation in MDE activities, paving the way for enhanced efficiency and effectiveness in software development. In addition to offering extraordinary automation capabilities for MDE, these emerging technologies create new opportunities for this paradigm to play new roles in development processes. One significant aspect to consider is the democratization of AI techniques. The power of abstraction and automation provided by MDE allows non-specialists to leverage AI capabilities seamlessly. This democratization can lead to more widespread adoption of advanced technologies and contribute to innovation across industries.
Moreover, MDE’s potential extends beyond traditional software development. It can be leveraged to build complex and hybrid systems that integrate physical, biological, social, and software components. Such systems are becoming increasingly crucial in modern societies, where interconnectedness and integration are paramount.